### v1

# Blockchain Economic Denial of Sustainability Attack: Exploiting Latency Optimization in Ethereum Transaction Forwarding

Abstract
Strategies related to the blockchain concept of Extractable Value (MEV/BEV), such as arbitrage, front- or back-running create an economic incentive for network nodes to reduce latency, including minimizing transaction validation time—a core feature to secure blockchain networks. A modified node, that neglects to filter invalid transactions in the Ethereum P2P network, introduces novel attack vectors. In this work, we formalize and evaluate a Blockchain Economic Denial of Sustainability (EDoS) attack, which can cause financial losses in traffic costs for operators of modified nodes. We 1) mathematically define the attack model, 2) identify thousands of empirical instances of this similar attack in the wild, 3) empirically measure the model parameters from our two monitoring nodes, and 4) conduct attack simulations on the local network to compare its performance with existing Denial-of-Service attacks. We show that an attacker can amplify network traffic at modified nodes by a factor of 3,600, and cause economic damages 13,800 times greater than the amount needed to carry out the attack. Despite these risks, aggressive latency reduction may still be profitable enough to justify the existence of modified nodes. To assess this trade-off, we 1) simulate the transaction validation process in the local network and 2) empirically measure the latency reduction by deploying our modified node in the Ethereum testnet. We conclude with a cost-benefit analysis of skipping validation and provide mitigation strategies against this attack.
1 Introduction
The journey of transactions in blockchain P2P networks en route to their ultimate validation remains understudied. The specifics of handling pending transactions vary according to the client’s software version and its unique implementation. Under the concept of Extractable Value (MEV/BEV) [9, 44]—obtaining extra profits by reordering transactions, centralized business entities have the incentive to reduce latency to deliver pending transactions. Being the fastest in this space (as little as 1 millisecond) attracts users, bots, and validators who profit from arbitrage or front/back running [9, 45]. This scenario bears a resemblance to the traditional stock market where trading firms physically shortened the network paths to stock exchange matching engines and optimized networking configuration for quicker trade execution [33]. Recently, we notice certain modified nodes tailor their configurations, sidestep transaction validation processes to shorten latency, and introduce invalid transactions that are not supposed to persist in the network.
Figure 1: Overview of Blockchain EDoS attack, illustrating how one invalid transaction amplifies in the network.
We formalize the Blockchain Economic Denial of Sustainability (EDoS) attack, a superset of previously proposed DoS (Denial of Service) attacks [35, 54]. In Figure 1, the adversary sends an invalid transaction (red arrow) to the modified node(s) that collectively propagate an invalid transaction (blue arrows) to the rest of the network, causing increased outgoing network traffic, and as a result, financial losses for those modified nodes. This paper fully describes such potential attack opportunities, documents the prevalence of similar attacks in the wild, empirically/simulatively models the network damage, and quantifies the trade-offs associated with skipping validations.
We first formally define the Blockchain EDoS attack and pinpoint the parameters that affect how an attacker’s invalid transactions can get amplified in the network. While considering the specifics of each software/version on transaction forwarding, our model establishes a framework to quantify the increase in both the amount and the economic cost of egress traffic at the modified nodes.
We next find that some centralized services propagate a significantly higher number of invalid transactions by analyzing the set of pending transactions (called “txpool”) for each entity. We identify 2,591 similar attack instances across 345 Ethereum blockchain addresses that target these services; this substantiates the attack feasibility. We classify those instances based on the attacker’s observed strategies and estimate the attack cost, size, and intensity.
To correctly estimate the overall impact of the attack and understand the topology of the network, we design two customized monitoring Geth nodes (the most popular Ethereum node client) to scan the activities on the P2P layer and record specific types of messages before they undergo processing, resulting 2.5 billion observations over 5 months. We formalize a new method for inferring the number of active peer connections, which serves as a cost-effective/ethical alternative to previously proposed methods [13, 10, 4, 34]. Furthermore, we show that 1.5% of nodes in our dataset exhibit lenient transaction validation. Many of those share an identical git commit, whose hash we do not find in any public software repository, implying those nodes come from the same entity.
Given the model we define and the parameters we obtain from our measurements, we show that the attacker can amplify outgoing network traffic at modified nodes by a factor of at least 3,600, and cause economic damages 13,800 times greater than the amount needed to carry out the attack.
In addition, we conduct attack simulations on the local P2P network and confirm that our EDoS attack can evict as many honest transactions (from both the txpool/the block) as the existing DoS attacks but at significantly lower costs.
Furthermore, we experiment to quantify the benefits of skipping transaction validation, specifically the amount of time saved for each validation process. We first fork our archive node (the node stores all the historical states of the blockchain) to simulate/dissect the process of Geth’s transaction validation in the locally controlled environment. We find that checking account status (nonce/balance) is the most time-consuming (86%). Moreover, we empirically measure the latency reduction by deploying the modified/regular nodes in the Ethereum testnet for 14 weeks.
Finally, we perform a benefit-cost analysis of a modified node based on time-to-money conversion [51, 45] and offer three possible mitigation tailored to Ethereum: 1) enforcing a stricter txpool policy, 2) postponing the validation process, and 3) introducing a node reputation system.
The contributions of this paper are as follows.
  • We formally define the Blockchain EDoS attack and show feasibility based on similar attacks in the wild.
  • By designing and deploying our custom monitoring nodes, we infer the Ethereum P2P network topology and calculate the attack amplification factors.
  • Simulations on the local network confirm the superiority of our attack over the existing DoS attacks.
  • We simulate, dissect and empirically measure the transaction validation process for latency reduction.
2 Background
We next present relevant background on blockchain peer-to-peer networks, and associated security issues.
2.1 Blockchain and P2P network
Three aspects of the blockchain ecosystem are particularly pertinent to our paper: 1) the formation of a P2P network by nodes, 2) the validations of transactions, and 3) the interaction with users outside the P2P network.
Ethereum nodes use the Kademlia [38] distributed hash table for connectivity. When a node launches, it initially connects to hardcoded bootstrap nodes, which provide a list of potential peers. The node identifier, known as “enode,” comprises the node ID (randomly generated from the node’s public key), IP address, and port number. The node calculates the distance between possible peers based on node IDs, divides them into buckets, and populates the list of peers, [^1] The node calculates the XOR of two node IDs: Keccak256(node1’s id) XOR Keccak256(node2’s id) as a distance, and group the possible peers into 256 buckets: for each $i$ (bucket), where $2^{i}\leq distance<2^{i+1}$) making peer selection arbitrary [16]. Geth by default allows 50 connections, while Erigon and Nethermind permit 100. Geth allocates one-third of connections to active peer searching and uses the other two-thirds for passively accepting inbound connections. Ethereum’s mainnet shares the same underlying P2P network with other chains (e.g., Ethereum Classic or Testnets such as Holesky). Geth nodes immediately disconnect from peers on different chains.
After establishing a connection, nodes initiate the exchange of messages to synchronize blockchain information (blocks and transactions), following the Ethereum Wire Protocol [15]. For every transaction, a node must validate the integrity of each transaction, which contains several parameters, regardless of encoding type. Ethereum is an account-based ledger, meaning each address maintains the balance and the transaction index called a “nonce” (starting from 0, incrementing per transaction). The amount of “gas” represents the computational costs of the transaction. We multiply by the gas price to obtain the total transaction fees. The transaction is either to 1) send “value” to another address, or 2) execute a smart contract where “data field” is used to specify the function and its input. The node checks the following conditions for each transaction:
[…]
  • The sender’s nonce is equal to/larger than the current nonce (i.e., greater than the past nonce) [^2] The acceptable nonce gap depends on each client’s implementation.
  • The transaction size is below a pre-defined limit (e.g., 128KB).
Beyond protocol-level verifications, each software incorporates supplementary checks to mitigate DoS (Denial-of-Service) attacks. For example, Geth employs the “gas bump rule,” requiring the gas fee for a transaction with the same sender and nonce to be higher than the original transaction with a specified increase (e.g., 10% in Geth). The node should not accept replicated transactions with the same address, nonce, and gas price. Li et al. [34, §5.1] or Yaish et al.  [54, §4.1] provide detailed explanations.
If the transaction is valid, it gets added to the buffer referred to as “txpool” (or mempool, tx-queue); otherwise, it should be dropped. Each software implementation imposes its own limit on the number of transactions accepted in the txpool; legitimate transactions may be discarded if the txpool reaches its maximum capacity.
When a new transaction enters the node’s txpool, the node broadcasts it to the rest of the network using two message options: 1) broadcast Transactions (0x02) – RLP (Recursive-Length Prefix)-encoded raw transactions that include all parameters, or 2) announcement NewPooledTransactionHashes (0x08) – transmitting only transaction hashes. Typically, the node broadcasts Transactions (0x02) to a small subset of nodes for efficiency, and announces NewPooledTransactionHashes (0x08) to the remaining nodes. If the node lacks information about a transaction based on its hash (0x08), it requests a peer to send transaction content via GetPooledTransactions (0x09), and the recipient node responds with PooledTransactions (0x0a). We describe software-specific implementations in §4.3.
A validator selects the set of transactions from its txpool and assembles the block. However, some transactions, despite passing the pre-check and being included in the txpool, never make it to the chain; these are referred to as dropped transactions. If the gas price is insufficient, validators may overlook these transactions. Subsequently, senders update the gas price, replacing the old transactions, which results in dropped transactions. In this paper, we assume transactions are “dropped” if they fail to be included in the blockchain within a span of more than seven days from their initial appearance in txpool. In Appendix A, we experiment to validate the use of a 7-day blockchain lookup. Some transactions are sent with insufficient funds or past nonce, making them ineligible for inclusion in the chain. We specifically use the term invalid transaction for those.
When users fetch the blockchain data or initiate a transaction, they typically interact with a P2P node through Remote Procedure Call (RPC) requests, and nodes execute these requests. Various publicly accessible RPC endpoints (e.g., Infura) allow users (rate limited) free access.
Txpool view providers (such as bloXroute, Eden, Chainbound) offer a “better” view of txpool in terms of latency. For example, bloXroute employs a Blockchain Distribution Network (BDN) infrastructure to improve blockchain scalability [31, 3]. The globally distributed nodes index transactions to reduce the size, internally propagate without validations (“cut-through routing”), and employ dynamic routing to overcome network overhead.
2.2 Security in P2P network
While a P2P network serves as the foundation of decentralized blockchains, sharing the same information among all peers is challenging, and often results in network forks [12]. Forks expose the network to double-spending [29, 22], eclipse [27] or selfish mining [19] attacks that all partition the network to block information flow. In the PoW blockchain, Luu et al. [36] proposed the “Verifier’s dilemma,” wherein rational miners opt to forgo block validation to allocate more time to PoW mining. Das et al.[11] demonstrated that PoW miners who bypass validation stand a significantly higher chance of winning the block compared to their mining power. The same applies to transaction latency especially if the blockchain supports decentralized exchanges, which allow transaction re-ordering attacks [9, 44]. As long as there is an economic incentive to capture transactions fast, centralized entities strive to minimize latency, which might make the P2P network more centralized and introduce new security vulnerabilities.
In this paper, we introduce an Economic Denial of Sustainability (EDoS) attack in blockchain P2P networks. Hoff [28] first presented the EDoS concept in a 2008 blog post. EDoS, akin to a traditional Denial of Service (DoS) attack, focuses on inflicting financial losses by targeting the victim’s traffic usage [52]. Attackers exploit the knowledge of a victim server hosted on a cloud service by manipulating traffic usage to inflate bills, making the victim’s operation economically unsustainable. The same issue applies to blockchain P2P networks where many nodes have public IP addresses, and use cloud platforms such as AWS.
3 Data
We use publicly available data (RPC nodes, txpool providers) from Flashbots [25], as well as measurement through our nodes deployed in a P2P network (Figure 2).
3.1 Public txpool data
Each node has a different view of the txpool, meaning that they hold a different set of pending transactions. A node that opens up its API connection to the public makes its txpool accessible (newPendingTransactions). Flashbots [^3] https://www.flashbots.net/ has been providing publicly available txpool information on “Mempool Dumpster” [25] every day, and we use its data from September 1st, 2023 to January 11th, 2024. Flashbots collects a set of pending transactions from 1) RPC providers and generic nodes – Infura, Alchemy, A-pool, [^4] A regular network node with the optimized peering setting whom Flashbots listens to. Flashbots’ local node, Mempoolguru [1], and 2) the infrastructure txpool providers – bloXroute, Chainbound, Eden. All the transactions contain transaction hash, source (i.e., which service/node), and observation timestamp (in milliseconds). Some portions of transactions include the details of transactions (e.g., sender, receiver, gas, value, and data fields). We remove repetitions (identical transactions) from the same source to avoid double-counting. Consequently, our summary statistics slightly diverge from Flashbots’.
3.2 Private monitoring node data
To effectively profile each node within a network, we gather peer information from our local Ethereum nodes. This is done by subscribing to the admin_peerEvents API, [^5] https://geth.ethereum.org/docs/interacting-with-geth/rpc/ns-admin which provides details such as node names, software, network IDs, compatible protocols, and IP addresses. This API alerts peer events to the subscriber in real-time, capturing both the addition and removal of peers, including those that connect/disconnect briefly. This information helps us calculate the number of active peer connections at each time, and better infer peer stability. For IP addresses, we call the ipinfo.io API [^6] https://ipinfo.io/developers/responses to obtain additional information such as Internet Service Provider (ISP) or Autonomous System (AS), hostname, timezone, region, city/country, and registered location. Besides regular full nodes, we also use our archive node to access the full state of the blockchain. In particular, we 1) retrieve the account balance/nonce at specific times (§5.3) and 2) fork the archive node to simulate the transaction validation process (§8.1). To accurately assess the attack cost amid price fluctuations, we obtain historical Ethereum/USD exchange rates through the Coingecko API. [^7] https://www.coingecko.com/api/documentation
Figure 2: An overview of our monitoring node.
We modify Geth to store all messages received on the P2P network layer between September 1st, 2023 and January 25th, 2024. We deploy two nodes in the network to 1) avoid a single point of failure and 2) increase the coverage and robustness of our estimates. Those monitoring nodes enable us to observe transactions before they enter the txpool as described in Figure 2; modified components are highlighted in yellow, whereas the standard Geth components are in grey. We have integrated a custom message dispatcher that filters and timestamps specific transaction propagation messages (e.g., 0x02 and 0x08 messages). These messages are subsequently forwarded to a message queue, then batch-processed by a goroutine worker for database entry. This configuration ensures there is no delay between the capture of a message and its logging in the database. For each transaction hash, we only record the first message to reduce the size of the dataset. We store the type of messages, the content of the transactions, the origin node IDs, and the timestamp. This is critical for 1) estimating the active peer connections (§6.1), and 2) identifying spamming behavior and modified nodes within the P2P network (§6.2). Our two Geth nodes operate on Ubuntu 20.04.2 LTS systems, one with an AMD Ryzen Threadripper 3990X (64-core, 2.9 GHz) and the other with a 13th Gen Intel(R) Core(TM) i9-13900KF. Both machines have 256 GB and 64 GB of RAM, respectively, and are supported by NVMe SSDs. We raise the limit of peer connections in both Geth clients to 1,000. Both nodes are located in Europe and are operated independently. We keep the peering default configuration (no change in whitelist/blacklist). In total, we capture about 2.5 billion (2,493,695,017) unique 0x02 and 0x08 messages from 36,815 peers in the aforementioned timeframe.
3.3 Ethics
Following Tang et al. [48] on ethics of measuring blockchain P2P networks, our work consists of passive measurements (i.e., not submitting transactions). We do collect the Ethereum node IP addresses, but this is already publicly available information. As a part of responsible disclosure, we contacted the services that we discovered are running modified nodes and could be victims of this attack.
4 Modeling EDoS attack
We turn to formally modeling the Economic Denial of Sustainability (EDoS) attack in detail.
4.1 Threat model
We first formalize an amplified EDoS attack on the P2P blockchain network (Figure 1). The attack leads modified nodes (i.e., victims) to forward invalid transactions, and increases their outgoing traffic usage, thereby causing financial loss. The attacker attempts 1) to cause economic damage to modified nodes, 2) to degrade the quality of the service provided by modified nodes, and 3) to disrupt high-frequency MEV bots or block builders that listen to the services provided by modified nodes. Beyond blockchain ecosystem participants, cloud service providers also potentially have an incentive to support attacks on modified nodes to increase traffic usage and, consequently, revenue.
We next delineate the attack process. Let $\mathcal{A}$ denote the attacker (node), $\mathcal{M}$ be a modified node, and $\mathcal{R}$ be a regular (unmodified) node. $\mathcal{A}$ follows the procedure in Figure 3 to produce an invalid transaction, $tx_{invalid}$, of size $a$ bytes.
Figure 3: Invalid transaction creation
  1. 1.
    If the target $\mathcal{M}$ does not check the account balance, $\mathcal{A}$ can generate different addresses and send an infinite number of invalid transactions. The transaction cost is zero since there is no chance of inclusion.
  2. 2.
    If $\mathcal{M}$ checks the balance but not the nonce, $\mathcal{A}$ can send transactions with a previously used nonce, which also costs nothing. One necessary condition is for $\mathcal{A}$ to rely on an account with a long transaction history (i.e., having a high nonce allows $\mathcal{A}$ to reuse many past nonces).
  3. 3.
    If $\mathcal{M}$ checks balance/nonce but does not use the gas bump rule (§2), $\mathcal{A}$ can replicate transactions by altering a parameter (e.g., a data field), while maintaining the same sender, nonce, and gas price. $\mathcal{A}$ could incur transaction costs if the original (or one of the duplicated ones) ends up on the chain.
In the first two cases, $\mathcal{A}$ can specify an unreasonably high gas price to prioritize their transactions, without additional costs. $\mathcal{A}$ can also combine multiple strategies (e.g., duplicating many invalid transactions using a past nonce).
Next, $\mathcal{A}$ selects its target $\mathcal{M}$. If $\mathcal{A}$ receives any invalid transactions from a given peer, that peer is most likely $\mathcal{M}$. Alternatively, $\mathcal{A}$ can monitor various centralized services’ txpools, identify the service(s) that accept(s) invalid transactions, and send $tx_{invalid}$ through RPC endpoints. $\mathcal{A}$ finally sends $tx_{invalid}$ to $\mathcal{M}$, which accepts/inserts $tx_{invalid}$ to its txpool and forwards it with the rest of the network, including other $\mathcal{M}$ nodes and $\mathcal{R}$ nodes.
When all $\mathcal{M}$ nodes are densely connected and run by the same entity (as empirically observed in §6.2), $tx_{invalid}$ reaches all $\mathcal{M}$ nodes. Even if some $\mathcal{M}$ nodes belong to multiple entities, $tx_{invalid}$ would reach out to the rest as long as one of $\mathcal{M}$ from each entity connects.
Regular $\mathcal{R}$ nodes keep receiving $tx_{invalid}$ as a new transaction because $tx_{invalid}$ never gets into their txpool.
4.2 Amplification factors
We want to precisely estimate the amplification factor to show how the attack scales up given the attacker’s input. The DoS/EDoS literature uses two metrics to quantify attack effectiveness: Traffic Amplification Factor (TAF) [32] and Economic Amplification Factor (EAF) [52], defined as
$TAF=\frac{B_{victim}}{B_{attacker}},\quad EAF=\frac{\lambda_{victim}}{\lambda_{attacker}}\ ,$
where $B_{victim}$, $B_{attacker}$ is the traffic from the victim and attacker node, respectively. In our scenario, TAF is the ratio of outgoing traffic generated by the attacker to the modified nodes, which corresponds to the original red arrow and the sum of the blue arrows in Figure 1, respectively. We calculate TAF for the network but not for a single modified node. $\lambda_{victim}$ (resp. $\lambda_{attacker}$) is the outbound traffic cost that the victim (resp. attacker) node pays for providers: in our case, all the modified nodes (resp. a single attacker). We can also re-write EAF as
$EAF=\frac{B_{victim}\cdot p_{victim}}{B_{attacker}\cdot p_{attacker}}=TAF\cdot\frac{p_{victim}}{p_{attacker}}\ ,$ (1)
where $p$ is the price per outgoing traffic cost based on the “pay-as-you-go” policy of cloud services. Typically, it should be easier for the attacker than for the victim to minimize their costs. Given that the attacker knows the IP address (i.e., location) of the modified nodes, they can launch the attack in the same data center and reduce their cost. This discrepancy (“price multiplier”) amplifies EAF.
4.3 Network waste
To estimate how one invalid transaction amplifies in the network, we first derive 1) the number of bytes a regular node $i$ receives based on the number of peer connections (see blue arrows directed toward each regular node in Figure 1), 2) the number of bytes the entire network wastes based on the distributions of active connections on each peer (all the blue arrows). This calculation only consider the public nodes in a P2P network, but not the users (MEV bots and block miners/builders) who 1) privately connect to txpool providers or 2) listen to transactions through RPC endpoints outside the P2P network. The actual amplification factor would be the combination of 1) the amount of traffic that goes to users who listen to modified nodes and 2) TAF. We indeed notice that some modified nodes shut down the outbound connections to reduce outbound costs. Yet, they still incur the cost of propagating invalid transactions to users.
A regular node only hears $tx_{invalid}$ from modified nodes (not from other regular nodes). This allows us to focus solely on the connections from modified nodes to each regular node. Let the number of active connections for each node $i$ be $x_{i}$. The number of the modified node connection is $x_{i}\gamma$ where $\gamma$ is the ratio of modified nodes. This is based on the assumption that the node discovery process is random, thus the number of connected modified nodes increases linearly along with the number of connections by expectation. As described in §2, Ethereum nodes broadcast the transactions in two ways: 1) 0x02 message called “broadcast” ($a$ bytes), and 2) 0x08 message called “announcement” ($32$ bytes). Although Ethereum Wire protocol [15] suggests propagating 0x02 to a small number of peers for network efficiency, there is no clear consensus on the implementation, leading to variations across different software clients.
For Geth, the propagation policy is consistent across versions; the node propagates the broadcast message (0x02) to a square root of connected nodes ($\sqrt{x_{i}}$) and just sends the announcement message (0x08) to the rest ($x_{i}-\sqrt{x_{i}}$)) defined in the function BroadcastTransactions. We call this the “square root policy.” The node picks up the subset of nodes uniformly randomly for every transaction.
In Erigon, just like Geth, the node broadcasts (0x02) a square root of connected nodes, and sends an announcement (0x08) to all peers (not the rest of the peers). While Erigon adopted a different propagation strategy after v2.49.0, we only focus on the square root policy given the small number of nodes after v2.49.0. The description of the change in Erigon propagation strategies is in Appendix B.
[…]
$f(x_{i})=\begin{cases}(0.14a+27.52)\gamma x_{i}&\pi=\text{``sqrt''}\ ,\\ a\gamma x_{i}&\pi=\text{``aggressive''}.\end{cases}$
[…]
(a) modified node ratio
(b) Propagation policy
Figure 4: Number of bytes a regular node $i$ receives based on the number of active connections.
As we already determine the waste for each regular node with a varying number of active connections, we proceed to calculate the total waste in the network by considering the distribution of peer connections, $g(x)$, which we empirically estimate in §6.1. There are $(1-\gamma)\cdot N=N_{\mathcal{R}}$ regular nodes where $N$ is the total number of nodes in the network. We multiply the number of regular nodes in the network, the distributions of peer connections $g(x)$, and the amount of waste for each node ($f(x)$) to get the overall network waste:
[…]
where $\int_{0}^{\infty}f(x)g(x)$ represents the expected waste for one average regular node. In practice, we integrate until 1,000 connections (to ensure convergence and exclude an impractically large number of connections) using Scipy. We finally divide this by the transaction size ($a$) to get TAF. An adversary could set up modified nodes on their own to increase $\gamma$, but these nodes would also be attacked, making this strategy unproductive.
5 Empirical examples of attack
This section identifies the similar attack instances seen in the wild, and characterizes the size, intensity, cost, and strategies, and actors (the attacker – sender address, the victim – centralized services) involved in those attacks. While we cannot definitively pinpoint the motivations of those attackers, those attacks are executed in a manner consistent with our EDoS attack (see the discussion in Appendix C).
5.1 Victims
We first attempt to find the instances of the attack in the txpool of centralized services as they are often the target of the attack. We find that some services deliver transactions much faster than others, but also propagate a significant number of dropped transactions. By comparing their txpools (from Flashbots’ data), we can practically figure out how fast, and how accurate their services are. For the latency comparison, we use the timestamp recorded in Flashbots’ data. For each on-chain transaction, we check how late each source propagates from the first moment Flashbots sees the transactions from any source. For accuracy, we calculate the ratio of the number of dropped transactions to the number of total transactions received. These numbers convey the quality of their propagation flow. Figure 5 (left) illustrates the bi-modal representation of latency vs. accuracy except for Flashbots’ “local” node. Each data point represents one week of data for each service. The $x$-axis is the median latency from all the transactions, whereas the $y$-axis is the ratio of dropped transactions. Services such as bloXroute, Chainbound, and Eden invest in infrastructure to optimize their propagation flow, resulting in potential latency reduction. Moreover, Flashbots’ database connects to bloXroute and Eden through gRPC, which is faster than the web socket commonly used by others (“delivery latency”). Additionally, a geographical advantage exists for services located close to the database—e.g., Flashbots’ local node likely operates a physical proxy to its database. The figure also suggests that those “fast” services fail to filter out transactions that are not supposed to persist in the network.
Figure 5 (right) is a co-occurrence matrix where each cell indicates the number of dropped transactions common to both services (the $x$- and $y$-axis) normalized by each day. The diagonal is the number of dropped transactions to each source—e.g., bloXroute has around 60,000 new dropped transactions per day. bloXroute, Chainbound, and Eden collectively accept a significantly high number of dropped transactions that are not seen in any of the other txpools, which suggests that they may have employed different propagation strategies from others. Those results hypothesize that a part of the latency reduction might also come from transaction validation process, which motivates our simulations in §8.1.
Figure 5: Comparing the centralized services’ txpools (left: accuracy-latency plot, right: dropped transactions between sources).
5.2 Identifying attack instances
We find that some Ethereum accounts send an unexpectedly large number of transactions in a short range of time. We set two (conservative) thresholds to extrapolate automated non-human spamming behavior: if 1) we detect over 100 transactions from a single sender in any of the transaction pools within a single block interval (12 seconds) and 2) more than 95% of those transactions fail to be included in the chain. The second check is necessary to filter out a large entity (e.g., a popular gambling site) that submits many transactions at once without spamming intent.
Based on this heuristic, we find 2,591 instances from 345 Ethereum addresses from Flashbots data from September 1st, 2023 to January 11th, 2024. Figure 6 summarizes the number of attacks over time (on an hourly basis). There are more than 60 instances of attacks on November 26th between 7–8 AM (UTC). Many of them appear to send transactions from accounts with zero balances. We find similar cases on September 12th, 5 PM, October 7th, 5 PM, and November 20th, 9 PM. The number of attacks does not appear to be correlated with the Ether price fluctuation.
Figure 6: Number of attacks observed per hour, over time (September 1st, 2023–January 11th, 2024).
5.3 Characterizing attack behavior
We first check the three types of invalid transactions discussed in §4.1: 1) insufficient balance, 2) past nonce, and 3) duplicate transaction. To check the balance and nonce of the relevant accounts at the time of transaction submission, we use our archive node to fetch the state of the blockchain at the time of the timestamp. Given that there is a network delay between when the transaction enters a P2P network and when Flashbots records the timestamp, we look at the state of the account two blocks before the observation timestamp (i.e., at least 12 seconds in the past). Of 2,591 attack cases, we confirm that 536 instances (from 224 unique addresses) come from accounts that do not have a sufficient balance. Two services—bloXroute and Eden—have included those transactions during our observation window. In particular, some addresses do not have any transactions (send or receive) in their history whatsoever (i.e., zero Ether balance). In addition, three services (Chainbound, Eden, bloXroute) have accepted transactions that were using a past nonce (and thus, are invalid) during our measurement interval. Just like we identify invalid transactions from accounts with insufficient balance, we retrieve the last nonce of the corresponding address and identify 62 attacks from 27 unique addresses that use previously used nonce.
Attackers also duplicate transactions in order to generate many invalid transactions. We categorize our identified attack instances based on the transaction parameters the attackers manipulate the most. There are four main parameters the attacker calibrates; 1) the gas limit, 2) the nonce, 3) the data field, and 4) the value of the transfer. This does not necessarily mean that the attackers only change those specific parameters, but sometimes they simultaneously permute several sets of parameters to increase the invalid transaction set, including other variables such as the recipient’s addresses, chain ID, and access list (introduced in EIP-2930). Table I summarizes the key statistics such as the transaction costs, intensity, and the size of the attack for each category. We first define key metrics: the size, cost, and intensity of the attack.
Size: we count the average number of transactions we observe during the attack and calculate the average transaction size (an RLP-encoded raw transaction in bytes). Since we only consider the transactions that enter one of the txpools Flashbots listens to, the number of transactions generated/sent by the attacker should be strictly larger.
Transaction costs: we calculate the “cost” of the attack as the amount of transaction fee (i.e., the effective gas price multiplied by the amount of gas used) for the original (on-chain) transactions that the attacker duplicates. For better comparison, we fetch the price data of Ethereum at the time of the attack and convert the total amount of Ether to USD.
Intensity: we compute the attack intensity based on the median of the timestamp interval (i.e., the time lag between two consecutive transactions observed in the txpool).
About 63% of attack instances calibrate the gas limit parameter, but only two addresses are involved. Those two addresses manage to send 1,581 transactions with only 0.26 milliseconds between transactions on average. The largest attack carries 23MB of data between blocks (12 seconds). Since, on average, only one transaction lands on the chain, the transaction cost to the attacker is around 3 US dollars. Chainbound seems to be the main victim of this attack.
Changing nonce (future/past) is also a common practice for attackers to generate a large number of invalid transactions. 300 addresses do so over 784 cases. Each of these attacks carries 227 transactions on average. Given that almost none of these transactions end up being included in the chain, the cost is exceedingly small.
The data field/value appears to be more costly than other methods (more than 30 USD) and carries a smaller number of transactions with around 10 milliseconds of interval. However, more (2.71 on average) services constantly accept those invalid transactions.
To sum up, Table I illustrates each strategy’s pros and cons; the attacker can choose to manipulate different parameters depending on what the attacker wants to maximize/minimize: costs, size, intensity, and the number of victims (i.e., the centralized services in §5.1 that accept those invalid transactions).
TABLE I: Summary statistics for each attack type.
Type # of cases addr
# of txs
# of on-chain tx
Total cost
(avg. in USD)
Txs interval
(avg. in ms)
Txs size
(avg. in bytes)
# of victims
gas 1719 2 1581.00 0.92 3.26 0.26 1496.2 1.04
nonce 784 300 227.24 0.03 0.05 9.22 120.2 1.09
data 71 30 167.23 1.11 31.97 10.56 563.0 1.70
value 17 13 116.35 0.94 132.00 7.91 481.7 2.71
6 Estimating model parameters
We next derive different sets of model parameters, which we then use to calculate amplification factors.
6.1 Peer connection distribution: $g(x)$
We formalize our method for estimating the number of active connections for each peer, $x_{i}$, and empirically computing its distribution, $g(x)$. Although studies on blockchain P2P network topology exist, none of the approaches directly translate to our context. The key novelty in our approach is to estimate active peer connections on the Ethereum mainnet without submitting test transactions (i.e., mediating ethical concerns and research costs). We use a 5-month dataset, encompassing 2.5 billion messages.
As outlined in §4, a unmodified Geth node $i$ has a number of peer connections $x_{i}$ and uniformly randomly chooses $\sqrt{x_{i}}$ nodes to broadcast (0x02) transactions, while sending announces (0x08) to the rest. Our node receives a 0x02 message with probability $\theta_{i}=\frac{\sqrt{x_{i}}}{x_{i}}$, which means the data distribution follows a binomial distribution with parameters $\theta_{i}$ and $m$, where $m$ is the number of messages we receive from each peer. Specifically, $m=m_{2}+m_{8}$ where $m_{2}$ and $m_{8}$ are the number of 0x02 and 0x08 messages sent by peer $i$, respectively. The probability (likelihood function) of the binomial distribution is defined as
$l(\theta_{i},m_{2},m_{2}+m_{8})={m_{2}+m_{8}\choose m_{2}}\theta_{i}^{m_{2}}(1-\theta_{i})^{m_{8}}\ .$ (3)
We derive an unbiased MLE (Maximum Likelihood) estimate by taking the derivative of the log-likelihood function, and reconstructing the number of peers $x_{i}$ as:
$\hat{\theta}_{i}=\frac{m_{2}}{m_{2}+m_{8}}\Leftrightarrow\hat{x}_{i}=\left(\frac{1}{\hat{\theta}_{i}}\right)^{2}=\left(\frac{m_{2}+m_{8}}{m_{2}}\right)^{2}\ .$ (4)
Since the estimate $\hat{x}_{i}$ becomes unreliable with smaller samples $m$, we also derive the variance of $\hat{\theta}$ and use it to exclude unreliable estimates on $x_{i}$. This is likely to exclude nodes with many peer connections, which require a larger sample to reduce the variance; thus we might underestimate the expected value of $g(x)$. However, 1) the percentage of nodes with large connections is disproportionately small, 2) the exclusion process does not lead to overestimating amplification factors, and 3) our large dataset helps alleviate this shortcoming. The details of how we derive the equations above and exclude unreliable estimates are in Appendix D.
We apply Eqn. (4) for peers (of our nodes) that 1) use Geth or Erigon (before v2.49.0), and 2) we exclude nodes with a non-public git commit string, which is a sign of software modification as we discuss in the next section. If both of our monitoring nodes observe the same peer, we take the average between two. In total, we manage to estimate the number of peer connections for 6,005 nodes: mean and median of 41 and 31, respectively. For those, each peer has sent an average (median) of 293,121 (81,327) transaction messages. To make the distribution generalized/smooth and get $g(x)$, we apply the kernel density function (KDE) based on SciPy’s implementation & Scotts’ method for the bandwidth selection. Figure 7 shows the resulting PDF (Probability Density Function) and CDF (Cumulative Density Function) for our estimate of active peer connections for each node based on data collected from our two measurement nodes. The strong alignment between the estimations between two nodes indicates our method is robust. As an additional check, we ask one Ethereum node operator to provide us with the node name, node ID, and the maximum number of peers. The node uses an unmodified Geth node and sets the maximum number of peers to 40. Our estimator tells us that this node is expected to have $33.68$ peers. Given that not all the nodes are stable over time, our estimate appears to be consistent with this piece of ground truth.
Figure 7: Number of active peer connections for each regular node: left (PDF with a KDE smoothing), right (CDF).
6.2 Ratio of modified nodes: $\gamma$
We estimate the ratio of the modified nodes ($\gamma$) from the set of peers we connect. As described in §3, we capture transactions before entering the txpool, so we can empirically label the modified nodes that forward invalid transactions to us. First, we prepare three sets of invalid transactions as introduced in §4.1 and §5 to examine peers’ validation process: 1) insufficient balance, 2) past nonce, and 3) duplication (gas bump rule). We first label peers that have ever sent invalid transactions of the first two kinds: insufficient balance and past nonce. Our approach may mislabel modified nodes if their blockchain status is unsynced. However, the fact that the node forwards invalid transactions still allows attackers to conduct the same attack on unsynchronized nodes.
With regard to duplications (which are state-independent), we flag transactions as “spamming” (i.e., duplicated) transactions when we see the attacker changing the data field to produce more than 10 transactions (e.g., just tweaking 1 byte) while keeping other parameters constant. We then label peers that forward duplicated transactions within one second, which indicates that the duplicated transactions stay in each peer’s txpool at the same time. That fact suggests the absence of a gas bump rule (or that the threshold was set to zero).
Surprisingly to us, we do not find many nodes that propagate transactions without sufficient balances or with past nonces even though we observe many at the RPC endpoints. We conjecture that those nodes might have already adopted a technique similar to the second mitigation we propose in §9.2 or restricted the outbound traffic.
To derive the ratio of modified nodes in the duplication case, we look at the set of nodes in our active connections and check whether any of those are from modified nodes. Figure 8 shows the ratio of connections to modified nodes at each snapshot (hourly) between December 15, 2023, and January 10, 2024. Both nodes $A$ and $B$ maintain approximately 15 connections over time with modified nodes; dividing the number of total number of connections (roughly 1,000) at each time interval gives us an average ratio of about 1.5%. In total, we identify 175 unique modified nodes.
Based on the HEAD of the git commit in the node name (which is generated when compiling the source code), we determine whether the node uses the publicly available version of the software or the private forks of the repository. 174 out of 175 peers share an identical git commit hash, which is not present in the corresponding GitHub repository. (We fail to collect information for one remaining node.) The details of how we extract the git commit hash and use it to determine whether a node was customized can be found in Appendix E. Further, we find that these modified nodes apply an aggressive propagation strategy ($\pi$=“aggressive”) – i.e., they only broadcast (0x02), but never announce (0x08), further increasing amplification effects.
Figure 8: Ratio of modified peers in our nodes’ connections.
6.3 Amplification factor
Based on the estimated parameters in the above sections, we finally calculate the amplification factors. We derive TAF based on the transaction duplication case, but the same technique could apply to insufficient balance or past nonce cases. Table II summarizes the model parameters and the final results. From the top, we refer to the size of transaction $a=560$ (bytes) in §5, modified node ratio $\gamma=0.015$, and propagation policy $\pi$=“aggresssive” in §6.2. We set the number of peers $N$ to 6,000 based on the statistics reported by [18, 17] after excluding non-Ethereum nodes. We plug in those numbers to Eqn. (2) and estimate the total waste in the network to be 2,037,613 bytes (2.0 MB). Given that the transaction size ($a$) is 560, the amplification factor is 3,638.
TABLE II: Estimated parameters for the model.
Parameter Description Estimate Source
a Tx size 560 §5.3
$\gamma$ Ratio of $\mathcal{M}$ modified nodes 0.015 §6.2
$\pi$ Propagation policy “aggressive” §6.2
N Number of nodes 6000 [18, 17]
Per tx per regular / modified node 345 / 22,640
Entire network 2,037,613 Eqn. (2)
Amplification BAF 3638
AWS ratio 0.2 Appendix F
EAF 13827
To quantify the economic impact of network waste (EAF), we next calculate the financial loss of spreading invalid transactions by modified nodes. We observe that many centralized services deploy their nodes in the cloud service (Appendix F), so we calculate the data transfer cost based on the “pay-as-you-go” policy of the cloud services. We choose AWS pricing [^8] https://aws.amazon.com/ec2/pricing/on-demand/#Data_Transfer_within_the_same_AWS_Region as an example. AWS does not charge for inbound traffic but for outbound traffic (significantly less for the traffic within AWS). Most cloud services follow a similar standard. If we assume that the cost of inbound traffic is zero, the EDoS attack incurs costs only for 1) the attacker (to send invalid transactions) and 2) the modified nodes (to propagate invalid transactions to the rest).
We compute EAF (economic amplification factor) based on TAF and the traffic price ratio, given by Eqn. (1). If the attacker strategically deploys its node in the cloud service co-located with modified nodes, the cost is 0–20 USD/TB (AWS US East) including the case when the attacker sends transactions to multiple modified nodes across regions. On the one hand, the modified nodes are organically connected to the rest of the P2P network, with 20% of internal AWS traffic (Appendix F) and 80% of external traffic to the internet. AWS dynamically charges external traffic starting from 90 USD/TB up to 10TB; the price gets discounted after.99990 USD/TB until 10TB, 85 USD/TB for the next 40TB, 70 USD/TB for the next 100TB, and 50 USD greater than 150TB Due to the price discrepancies between the attacker and the modified nodes (i.e., price multiplier), EAF is 13,827 for the first 10GB of non-AWS traffic (12.5 GB for all). Figure 9 illustrates the decrease in EAF ($y$-axis) based on the amount of outgoing traffic ($x$-axis) due to AWS’s dynamic pricing. The attacker’s economic benefit marginally decreases over the scale of the attack.
Figure 9: EAF based on the traffic usage of a modified node.
7 Attack simulation in the local P2P network
We evaluate our EDoS attack in the local P2P network simulation and compare it with the existing DoS attacks.
We set up a local P2P network, in which we deploy our Geth node (modified or unmodified), attack our own nodes, and then assess both txpool congestion and transaction inclusion in the blocks. We base our attack on Yaish et al.’s publicly available repository, [^10] https://github.com/AvivYaish/SpeculativeDoS which allows for a direct comparison of our approach with the two existing DoS attacks, namely 1) Baseline, that is, sending valid transactions with higher gas prices (naive eviction strategy), and 2) MemPurge [54], that is, sending future latent (invalid) transactions by extending the DETER method [35]. In these two attacks, we do not modify our node.
We set up one validator, 80 honest accounts, and a varying number $x$ of malicious accounts, all of which interact with our node through the RPC endpoint. First, each honest account sends 64 transactions to collectively fill up the transaction pool with valid transactions. Each attacker then begins submitting attack transactions. For our EDoS attack, each attack account attempts to transfer an amount (1 ETH) higher than its current account balance, and starting from nonce 0, which is lower than the current nonce – i.e., these transactions are invalid. Each attack account then gradually increments the nonce to generate 32 invalid transactions (i.e., a total of $32x$ attack transactions) with the intention of evicting honest transactions.
We measure 1) the ratio of honest transactions in the txpool, 2) the ratio of honest transactions in the final block, and 3) the number of attack transactions in the block (i.e., the attack cost). We calibrate the number of malicious accounts controlled by the attacker to measure the changes in those three metrics. Figure 10 depicts the relationship between those three metrics ($y$-axis) and the number of malicious accounts controlled by the attacker ($x$-axis).
Figure 10 (left) illustrates that our EDoS attack can evict as many honest transactions as the Baseline attack. MemPurge requires even a larger number of attack accounts to replace honest transactions, as the node pushes back future transactions to the queue if the txpool is already full.
Figures 10 (middle and right) illustrate that our EDoS attack can exclude honest transactions not just in the txpool but block as well because the attacker’s invalid transactions occupy the txpool. Furthermore, these invalid transactions are never included in the block, resulting in zero attack transaction costs. MemPurge gradually displaces honest transactions in the block, while the attacker pays for one transaction per attack account. The Baseline attack replaces honest transactions with a smaller number of attack accounts. Yet, all the attack transactions are included in the block, resulting in a significantly high(er) attack cost.
In summary, our attack achieves the same level of txpool/block eviction rate as the two existing DoS attacks but with significantly lower (zero) attack transaction costs. This advantage allows the attacker to keep refreshing the txpool with its new invalid transactions, causing the modified node to continuously notify its neighbors, thereby incurring egress traffic loss—EDoS attack. In Appendix G, we describe more detailed attack scenarios for the EDoS attack and roughly estimate the maximum traffic monthly costs incurred at the modified node to be 88,904 USD per node (8M USD in aggregate) when an attacker saturates outgoing traffic with invalid transactions.
We open-source our attack code in our repository [^11] https://github.com/taro-tsuchiya/BlockchainEDoS and document the attack implementation details.
Figure 10: The ratio of honest transactions in the txpool (left), the ratio of honest transactions in the block (middle), the number of attack transactions in the block (right), over the number of attack accounts (x-axis).
8 Estimating transaction validations
In this section, we conduct two experiments: 1) simulation, and 2) empirical measurement, focused on the latency benefits to understand the economic rationale behind skipping certain transaction validation steps.
The Geth node performs three sets of validations: 1) state-independent validations, such as checking transaction size, parameter range, minimum gas price, and signature; 2) state-dependent validations, such as verifying account balance and nonce; and 3) txpool checks such as the existence of the replicated transactions (“gas bump rule” in §2). The Geth node executes the above three validations at the following three stages: Acccept, Foward, and Update.
Accept: Each node validates transactions received from its peers or RPC endpoint and places them into a “(non-executable) queue”: ValidateTransaction(), ValidateTransactionWithState(), Add(), enqueueTx().
Forward: the node re-validates/moves transactions into txpool, and starts forwarding them to peers: promoteExecutables(), promoteTx().
Update: Upon the arrival of a new block, the node validates and removes transactions that are already included in the blockchain or become invalid due to a change in the blockchain state: demoteUnexecutables().
8.1 Simulating validations
We first create an isolated fork of the Ethereum blockchain within our local network to estimate the Geth validation processing time. We use a modified Geth client, integrated with the Pebble database. [^12] Geth can run on either Pebble or LevelDB as its underlying database. This approach allows us to simulate a controlled environment without impacting the live P2P network. Our objective is to examine how different types of validations contribute to the overall processing time and identify potential areas for efficiency improvements. Specifically, we run two functions in the first “accept” stage: ValidateTransaction() and ValidateTransactionWithState() for 1,000 consecutive blocks, starting from block number 18,140,000, in a total of 136,437 transactions. Simulating the initial stage (“first”) would suffice to estimate validation time, as the node caches the results of the account fetch. We conduct our experiment on a Ubuntu 22.04.2 LTS server, equipped with an AMD Ryzen Threadripper 3990X 64-Core Processor and a Corsair MP400 NVMe disk.
Our experimental results indicate that running the first validation stage takes roughly 1 millisecond per transaction. In particular, state-independent validations constitute roughly $12\%$ of the total validation time, with an average duration of $0.08$ ms (std=$0.04$) per transaction while state-dependent validation requires $86\%$ (mean=$0.89$ ms, std=$1.19$) of the total time. Within the state-dependent validation, nonce checks were the most time-consuming, accounting for about 82.2% of the total validation time. This significant portion largely comes from 1) a large blockchain state size, 2) disk accesses, and 3) the inherent complexities of the Merkle Patricia Trie (MPT) structure used for managing the Ethereum blockchain state. Refer to the detailed analysis in Appendix H.
8.2 Empirically measuring validations
The previous experiment solely examines validation processing time, omitting txpool or the timing of transaction arrival. We next empirically measure the latency difference between modified and regular nodes. Since our modified node may forward attackers’ invalid transactions, we choose to run the node in the Ethereum testnet “Holesky” instead of the mainnet. Our modified node bypasses the first (“accept”) and the second (“forward”) validation stages, but keeps the third (“remove”) validation stage, because the node gets congested by past transactions and stops inserting new transactions. To validate our implementation, we send one test invalid transaction (insufficient balance and past nonce) through our node’s RPC and verify that this transaction appears in our modified nodes’ txpool. To limit confounding factors, we alternate between running an unmodified and a modified node on the same machine. This allows us to control for 1) the specs of the machine (Ubuntu 22.04.4 LTS, with an AMD EPYC 9124 16-Core processor) and 2) the network topology by connecting to a similar set of peers using the same public key. To further eliminate the effect of peer connections, we increase the peer limit from the default (50) to 300 to connect to as many active peers as possible (based on the observation by Zhao et al [55] on the Goerli testnet). We sync the blockchain ahead of time and run our node from April 4, 2024, 1 PM (UTC) to July 12, 2024, 1 PM. We switch the modified node and the regular node weekly or bi-weekly basis, resulting in 7 weeks for each node. We capture around 34 million transactions in total.
To compare performance, we monitor each txpool and record the timestamp when the node finishes processing/validating each transaction. We use a metric called “inclusion time” [50]: the timestamp of blockchain minus the first observation timestamp in our txpool. The longer the inclusion time, the faster the node discovers and processes transactions. Figure 11 illustrates the distribution (until 120 seconds for visibility) for inclusion time for the modified (right) and the regular (left) node, respectively. We exclude periods when the market behaves abnormally (e.g., due to attacks or testing); refer to Appendix I.
The median inclusion times are 10.44. and 10.62 seconds for the regular node and for the modified node, respectively. This result suggests the modified node delivers transactions faster than the regular node as expected.  [^13] We use the Mann–Whitney U test—non-parametric test robust to outliers, to confirm statistical significance. Despite the long-term measurement, market conditions vary significantly on a daily basis (especially in the test net), making it difficult for us to make a fair comparison between nodes. We discuss the limitations of this experiment in Appendix I.
Figure 11: Time inclusion for a modified/regular node
9 Discussion
We first discuss if, despite the potential for losses coming from attacks, it could still be economically rational for an operator to modify nodes. Second, we describe the limitations and the possible mitigations against the attack.
9.1 Cost-benefit analysis
We summarize the benefits and the costs for each player: an attacker node ($\mathcal{A}$), a modified node ($\mathcal{M}$), and a regular node ($\mathcal{R}$). The attacker can launch an EDoS attack to induce financial losses for modified nodes. For instance, an attacker economically competing with the entity operating modified nodes could use the attack to weaken the competition. The attacker can also use its invalid transactions to displace legitimate transactions, rendering the service unusable (i.e., causing a DoS) as shown in §7.
[…]
Based on §7 and §8.1, while omitting validation checks leads to a possible reduction in processing time by the order of milliseconds, the marginal time savings may not necessarily justify the potential damages/risks from a Blockchain EDoS attack (in terms of MEV profits). Modified nodes engaging in these “optimizations” also expect to face degradation in the quality of service provided, due to the increasing number of invalid transactions (as shown in §7). With this in mind, a rationally economic node would still continue skipping validations as long as the economic benefits (including other forms of payments such as user subscription fees) outweigh any financial losses incurred from the attack.
Regular nodes can expect to receive transactions faster by connecting to the modified nodes but receive more invalid (less valid) transactions, which consume network capacity and CPU, indicating that our attack poses a threat to the security of all the players in a P2P network. (Our attack does not cause congestion in these nodes’ txpool, since regular nodes discard invalid transactions.)
TABLE III: Estimation of profit increase vs. latency reduction per block in Ethereum based on Wahrstätter et al. [51]. Calculations assume an ETH price of \$2,500, and an average MEV bid of 0.06 ETH.
Submit the block at $0.409s$ $1s$ $2s$ $2.5s$
For Every 1ms Latency Reduction
% Profit Increase $0.034\%$ $0.031\%$ $0.018\%$ $0.0074\%$
Profit Increase (ETH) $\approx$ 0.000020 $\approx$ 0.000019 $\approx$ 0.000011 $\approx$ 0.0000044
Profit Increase (USD) $\approx$ \$0.050 $\approx$ \$0.047 $\approx$ \$0.028 $\approx$ \$0.011
9.2 Limitations and mitigations
A few key assumptions affect our estimation model. First, we assume that the set of nodes we are connected to is a representative sample of the overall population. However, if our node(s) is/are more likely to connect to the node with many open connections, we might overestimate the distribution of peer connections $g(x)$.
Second, while our method of estimating the number of active peers requires the assumption of the network being static over time, we mitigate this concern by only including the nodes with a significant number of messages—stable nodes, thereby reinforcing the validity of our assumption.
Third, while our model assumes that the node transmits each invalid transaction once, in practice, the node can resend the same transaction multiple times, which could significantly increase the amplification factor.
Possible mitigations. Previous work suggests a few countermeasures to prevent EDoS attack [5]: 1) testing the client (Turing test, cryptographic puzzle) to exclude non-human requests or 2) blocklisting users by detecting abnormal traffic. Those solutions do not directly translate to our context because 1) a blockchain node is not interactive (i.e., a human would not be able to answer challenges at the requested rate) and 2) node identity can be easily spoofed. We devise three solutions tailored to blockchain P2P networks.
First, one can enforce a stricter txpool policy (checking balance, nonce, gas price bump) on modified nodes. Upon recovering more stringent transaction checks in the modified nodes, attack vectors diminish. However, as long as profits could potentially exceed losses on modified nodes, we cannot expect the solution to be universally implemented.
Second, one could postpone the validation process by introducing a two-step propagation mechanism. The idea is similar to the solutions proposed by Das et al [11] for delaying block validation in the context of the Verifier’s Dilemma. After the modified node receives transactions, it 1) directly relays them to users without validation, ensuring latency reduction, 2) validates transactions afterward, and 3) sends them to the rest (or does not forward them at all). This two-step approach does not incur additional egress traffic. However, modified nodes still pay the cost of delivering invalid transactions to users with an unreliable txpool view, creating a risk for users to include invalid transactions.
Third, one could employ a reputation system on regular nodes (e.g., similar to ISP blocklisting [6]). If the (regular) nodes have a mechanism to disconnect from the modified nodes that propagate many invalid transactions, the modified nodes would stop hearing new transactions, getting isolated from the rest of the network. This approach might yield false positives (i.e., legitimate nodes being blocked due to discrepancies in synced statuses) or lead to a cat-and-mouse game (i.e., adversarial nodes could come up with techniques to bypass/game the reputation system) [23].
10 Related work
We discuss related work by examining previous efforts in P2P network measurements and network security.
10.1 Measurement on P2P network
Most of the measurements on blockchain P2P networks focus on describing the topology of the network such as network size [30, 43], influential nodes [34, 37], stability or longevity of the connections [42, 10], and latency between peers [12, 21, 47, 50]. Some websites summarize those statistics in real-time for various networks (e.g., Bitcoin [2, 7, 46]; Ethereum [18, 17, 39]; Monero [41])
In particular, there has been extensive work on identifying active peer connections to further understand the propagation flow over the network. The techniques developed include: utilizing latency between peers [43, 10], isolation property and orphan transactions [13], connection freshness [4], and gas fee change [34, 55]. However, these works are either 1) chain-specific and do not apply to Ethereum, and/or 2) involve active measurement (e.g., submitting transactions), which is expensive and could disrupt the main network.
Our proposed method in §6.1 works on the Ethereum main network 1) without intervening in ongoing activities and 2) without incurring any additional costs apart from collecting data. Our method only considers the number of connections and does not determine whether an active edge exists between two specific nodes, making it harder to compare its accuracy/performance with previous proposals.
Another body of measurement research focuses on characterizing nodes, e.g., centralized entities such as mining or relay nodes [40, 10, 34, 35], AS/cloud services [20, 21, 43], geo-distributions [42, 30, 10, 37, 4, 8], and peer technical specifications (i.e., bandwidth, CPUs) [47, 8]. Closest to our work is the literature about the client (software) types – for Bitcoin [42], Ethereum [30, 8, 24], and Zcash [10]. Those works mostly focus on the diversity of software and/or versions, with relatively scarce attention on the implementation and performance. To the best of our knowledge, there has not been any exploration of customized clients, a gap our study aims to bridge.
10.2 P2P network security: DoS and spam
Due to the rise of centralized services such as mining pools, exchanges, and relay nodes, Denial of Service (DoS) attacks on those services can have a disastrous impact on the network [49]. The most directly relevant work is by Li et al [35] (and its extension by Yaish et al [54]), which exploits the txpool’s propagation policy by sending transactions with a future nonce to evict legitimate transactions in the txpool. Further, Wang et al [53] developed a fuzzer to automatically discover txpool vulnerabilities. Zhou et al [56] document the spamming behavior of MEV bots as a computational overhead to the network. Our paper describes a new Blockchain DoS attack, exploiting the lack of transaction validations—a superset of the above DoS-related works, as shown in §7), and empirically derives the amplification factors. We also reference some metrics from non-blockchain DoS literature such as amplified DoS and EDoS attacks from Kumar [32] and Wang et al. [52], respectively.
11 Conclusion
A reduction in latency creates an economic value as a form of MEV/BEV, and some entities modify the software client (“modified node”) to shorten the transaction validation process. We formalize a Blockchain EDoS (Economic Denial of Sustainability) attack, identify similar attack instances in the wild and empirically measure and simulate the attack impact. Our model concludes that an attacker can amplify its original traffic by 3,600 times, and the financial loss by 13,800 times. A lack of transaction validation not only incurs a financial loss on the modified nodes but also causes an uptick in the number of invalid transactions in the network. This degrades the user experience, creates negative externalities by impacting other nodes, and poses a threat to the overall stability of the blockchain ecosystem.
Acknowledgments
This research was partially supported by Carnegie Mellon CyLab’s Secure Blockchain Initiative and by the Nakajima Foundation.
Appendix A Blockchain lookup
We systematically examine all transactions observed in Flashbots’ dataset to ascertain their eventual inclusion on the blockchain, differentiating them from dropped transactions–those that enter the txpool but fail to be included on the blockchain. Our analysis involves comparing the blockchain information up to the future seven days from the observation timestamp. There is a potential for mislabeling a transaction if it takes longer than seven days to be a part of the blockchain. To substantiate the adequacy of the 7-day lookup window, we conducted an experiment to check the variation in our results (i.e., optimizing the extent of blockchain lookup required to define dropped transactions). Utilizing 100 days of data starting from September 1st, we compute the number of dropped transactions for each day, calibrating the length of the blockchain lookup by extending the reference point into the future.
In Figure 12, the $x$-axis represents the number of future days of blockchain data utilized, while the $y$-axis denotes the level of dropped transactions (left: the mean dropped transactions per day, right: normalized by the amount of dropped transactions when $x=0$). The graph illustrates a marginal decrease in the number of dropped transactions as the window size increases. This trend implies that examining more than seven days of future blockchain is adequate for identifying dropped transactions. Specifically, extending the blockchain lookup by one additional week only results in a negligible 0.2% change in transactions. If a transaction persists in the network for over one week, chances of being incorporated into the blockchain become exceedingly low.
(a) Mean
(b) Normalized
Figure 12: Number of dropped transactions.
Appendix B Erigon propagation policy
This supplementary section illustrates the change in Erigon’s transaction propagation policy. Erigon had adopted “square root policy” just like Geth until v2.49.0. Refer to SendMessageToRandomPeers used in BroadcastPooledTxs. However, there have been two major changes to this initial implementation.
On August 23rd, 2023 (#8030), Erigon stopped the square root policy and started to broadcast (0x02) to every peer the node connects to. The intention is to propagate the block to every node, but the change in SendMessageToRandomPeers function affects the transaction propagation simultaneously (after version v2.49.0). Next, on Dec 4th, 2023 (#8271), Erigon decided to broadcast (0x02) to a constant number of 3 peers and announce (0x08) to 6 peers (after v2.55.0). The intention is to reduce the burden of outgoing traffic. eriogn’s analysis [14] illustrates that outbound traffic reduces from 5.5-6.5 MiB/s to 3-3.5 MiB/s. Our reconstruction method (§6.1) does not apply after v2.49.0 because an Erigon node 1) does not differentiate between 0x02 and 0x08 until v2.55.0, and 2) does not announce to all peers after v2.55.0.
Appendix C Manual analysis of the empirical attack
We manually look at some attackers’ addresses and list two possible motivations for the attack. The first interpretation is to disrupt the service or the network. If the account sends many transactions with insufficient addresses or past nonce, the attacker has no intention of making transactions on-chain, but solely to cause a disturbance to the service/network (i.e., EDoS attack).
The second interpretation is to increase the chance of transaction inclusion. The intention is different from our attack model. We look into two accounts (0x3d9e…, 0x443d…) that slightly alter the gas limit, and send thousands of invariant transactions in one block slot. Those accounts appear to be the arbitrage MEV bots. By sending thousands of transactions at once, the transactions may reach out to the nodes faster than others or fill up the txpool to prevent other transactions from being executed. This behavior is partly discussed in another study by Zhou et al [56]. While many of their on-chain transactions get reverted in the middle of the execution, there are some successful complete arbitrage transactions. One transaction (0x3d1d…) has made 1.42 Ether by taking an arbitrage between Uni Swap and Sushi Swap, which would be 2,909 USD at the time of execution. As long as the cost (i.e., the amount of gas fee that the MEV bot pays for on-chain transactions and the computational cost of generating invalid transactions) does not exceed the profit of successful arbitrage, the bot has an incentive to spam the network.
Despite different attack intentions (or scale), the attacker crafts many invalid transactions and disrupts the service/network. The prevalence of the attack in the current P2P network corroborates the practicality/feasibility of our EDoS attack.
Appendix D MLE estimate of peer connections $\hat{\theta}_{i}$ and $\hat{x}_{i}$
Figure 13: The relationship among $m$, $x_{i}$, and the amount of error.
This section complements the mathematical formulation presented in §6.1. From the likelihood defined in Eqn. (3), log-likelihood function is
$\ln l(\theta_{i},m_{2},m_{2}+m_{8})=\ln{m_{2}+m_{8}\choose m_{2}}+m_{2}\ln\theta_{i}-m_{8}\ln(1-\theta_{i})\ .$
We take the derivative of this function w.r.t $\theta_{i}$ and set the equation to 0 to maximize the likelihood, which produces the unbiased parameter $\hat{\theta}_{i}$ (i.e., the ratio of 0x02 messages).
$\frac{d\ln l(\theta_{i},m_{2},m_{2}+m_{8})}{d\theta_{i}}=\frac{m_{2}}{\theta_{i}}+\frac{m_{8}}{1-\theta_{i}}=0\Leftrightarrow\hat{\theta}_{i}=\frac{m_{2}}{m_{2}+m_{8}}\ .$
This allows reconstruction in Eqn.(4).
We next calculate the variance of $\hat{\theta}_{i}$ to quantify the level of uncertainty for the estimate. Assuming that the MLE estimator is asymptotically normally distributed (i.e., when the sample size is large enough), the estimator $\hat{\theta}_{i}$ is normally distributed. Based on the Fisher information, the variance ($Var[\hat{\theta}]$) can be derived by calculating the 2nd derivative on the log-likelihood w.r.t. $\theta_{i}$.
$\frac{d^{2}\ln l(\theta,m_{2},m_{2}+m_{8})}{d\theta_{i}^{2}}$
$=-\frac{m_{2}}{\theta^{2}}+\frac{m_{8}}{(1-\theta^{2}}\quad\text{Using $\hat{\theta}_{i}=\frac{m_{2}}{m_{2}+m_{8}}$}$
$=-\frac{\hat{\theta}_{i}(m_{2}+m_{8})}{\hat{\theta}_{i}^{2}}+\frac{(1-\hat{\theta}_{i})(m_{2}+m_{8})}{(1-\hat{\theta}_{i})^{2}}=-\frac{m}{\hat{\theta}_{i}(1-\hat{\theta}_{i})}\ .$
The variance is the negative inverse of the 2nd derivative; thus $Var[\hat{\theta}_{i}]=\frac{\hat{\theta}_{i}(1-\hat{\theta}_{i})}{m}$. It is at most (bounded by) $\frac{1}{2\sqrt{m}}$ when $\hat{\theta}=\frac{1}{2}$ (maximum). Within two standard deviation (std) $\sigma_{\hat{\theta}_{i}}$ (which covers around 95%), a confidence interval (CI) of $\hat{\theta}_{i}$ is $\hat{\theta}_{i}\pm 2\sigma_{\hat{\theta}_{i}}=\hat{\theta}_{i}+\frac{1}{\sqrt{m}}$. However, after the reconstruction of $\hat{x}_{i}$ using Eqn.(4), the variance of $\hat{x}_{i}$ does not appear to have a closed-form solution, so we use heuristics to determine the credibility of our estimate $\hat{x}_{i}$. In particular, since the variance of $\hat{x}_{i}$ would expand more quickly when $\hat{\theta}_{i}$ is small, we exclude the uncertain estimates based on the level of $\hat{\theta}_{i}$ and $Var[\hat{\theta}_{i}]$. We calculate $\hat{x}_{i}$ from two $\hat{\theta}$s: 1) $\hat{\theta}_{i}$, and 2) $\hat{\theta}_{i}+(1/\sqrt{m})$ (two std apart), and refer to the difference of the two points as an “error,” $\varepsilon$. If the error is more than 10 (i.e., the estimated number of peers $\hat{x}_{i}$ likely to deviate more than 10 peers), we exclude the estimate.
$\varepsilon=\underbrace{\left(\frac{1}{\hat{\theta}_{i}}\right)^{2}}_{\theta_{i}=\hat{\theta}_{i}}-\underbrace{\left(\frac{1}{\hat{\theta}_{i}+(1/\sqrt{m})}\right)^{2}}_{\theta_{i}=\hat{\theta}_{i}+(1/\sqrt{m})}=x_{i}-\left(\frac{1}{\sqrt{x}}+\frac{1}{\sqrt{m}}\right)^{-2}\ .$ (5)
We next look at the relationship between $x$, $m$, and the error $\varepsilon$. Figure 13 (left/middle) illustrates the change in the level of error by increasing $m$ and $x_{i}$. When the number of observations $m$ increases, the error marginally decreases. When the peer connection $x$ increases, the error slowly increases. Figure 13 (right) shows the relationship between $x$ and $m$ while fixing the amount of error. A much larger number of samples is necessary for the node with many connections to obtain a reliable estimate.
Appendix E Customized nodes
Figure 14: Number active peer connections in our nodes’ connections. The $y$-axis does not start at 0.
Figure 15: Number active customized peer connections in our nodes’ connections.
We find a cluster of nodes that customize the existing clients and collectively develop the software privately. As noted in §LABEL:subsec:other_data, we collect the node name to profile P2P network nodes. Node names (e.g., Geth/v1.13.4-stable-3f907d6a/linux-amd64/go1.21.3) generally follow a convention indicating 1) software name, 2) version/branch and the first 4 bytes of the last git commit hash, 3) operating system, and 4) programming language and its version. We use regular expressions to extract the version and git commit hash. If the node owner forks existing projects, modifies any part of the code, and updates the repository, a different git hash is generated. To determine whether the node is customized, we call the Github API [^16] https://docs.github.com/en/rest/commits/commits?apiVersion=2022-11-28 to check whether the commit publicly exists in the corresponding software repository. We call nodes with non-public git commit hashes customized nodes. Given the space of possible hashes (4 bytes: $16^{8}$), the probability of generating two identical hashes by mere coincidence is extremely low, which suggests the cluster of nodes being operated by the same entity.
We identify many clusters of customized nodes. For example, the largest cluster has 250 nodes, and all reside in the same ISP (data center) in one (US) city. There are also two clusters (200+ different IDs) that appear to be operated by the same entity, but those nodes are spread out across different ISPs and 19 cities. We infer that they may belong to one relay service that attempts to reduce the latency by deploying nodes globally. Figure 14 illustrates the number of active connections in our two deployed nodes (based on the timestamp of connection/disconnection for each peer), which shows the stability of our nodes over time. Figure 15 depicts the number of customized nodes; around 100 (10%) are customized nodes.
Appendix F Software client, ISP
We investigate software client market share to validate our choice of transaction propagation policies in §4.3. To exclude unstable/non-functioning nodes, we produce statistics from nodes that maintain peer connections to our two nodes. Our reported statistics are thus different from those at websites such as Etherscan [18], which include all peers from other chains in the P2P network. We observe a software market share that remains stable over time. Geth, Erigon, and Nethermind have 75%, 20%, and 5%, respectively. We can thus focus on Geth and Erigon in our model.
We next look at the number of nodes in each cloud service, which helps calculate the traffic costs of the attacker/modified nodes. Our data shows the concentration of nodes in a handful of autonomous systems (AS). Deploying the blockchain node in the cloud is expensive, but brings potential latency reduction benefits by co-locating with other nodes in the same data center. The top four ASs are all cloud services and account for more than 50% of the active nodes. In particular, 20% of all active nodes appear to be hosted on AWS.
Appendix G Attack scenario
We aim to estimate the highest possible traffic costs that modified nodes would incur when an attacker saturates their outgoing traffic with its invalid transactions. The attacker can utilize 400 Ethereum blockchain accounts (as shown in Figure 10) to evict existing transactions in the txpool and keep refreshing it with the new invalid ones, causing the modified nodes to perpetually forward invalid transactions to their neighbors. The attacker can also send transactions from multiple nodes with different IPs, and distribute the attack to multiple modified nodes to scale up the attack until the bandwidth limit (i.e., Distributed EDoS attack). This helps attackers circumvent any network management tools deployed by modified nodes (e.g., anomaly detection—one of the solutions to classical EDoS attacks). We partly reference Kumar’s work [32] to calculate the total link capacity of intermediate nodes (i.e., modified nodes in our case). Table IV summarizes the assumptions and the final cost of modified nodes. We use the bandwidth of AWS EC2 instances for attacker/modified nodes and the minimum bandwidth requirement by Geth for the regular nodes. [^17] https://geth.ethereum.org/docs/getting-started/hardware-requirements In our case, the scale of the attack (the amount of waste) is mainly constrained by the number/bandwidth of modified/regular nodes since the attacker’s traffic gets easily amplified by TAF. Based on our calculation, each modified node could spend at most 88,904 USD, or roughly 8M USD per month in aggregate if all the modified nodes are from the same entity.
Alternatively, the attacker might opt to perform this attack moderately over an extended duration to evade detection, rather than maximizing traffic volume over a brief period. Previous literature [5] highlights the potential “stealthy nature” of the EDoS attack (compared to DoS) where the attacker carefully chooses the level of traffic under the detection threshold set by the victim. It might be more economically effective to control the level of the attack given the marginal decrease in EAF (Figure 9).
TABLE IV: One attack scenario ($\mathcal{M}$: modified node, $\mathcal{R}$: regular node) with link capacity constraints.
Capacity per node # of nodes Capacity (network) Cost
$\mathcal{M}$ $\beta_{\mathcal{M},out}$ $\gamma N$ $\beta_{\mathcal{M},out}\gamma N$
2.5 GB/second 90 225 GB/second \$89K/node
6480 TB/month 583,200 TB/month \$8M/all
$\mathcal{R}$ $\beta_{\mathcal{R},in}$ $(1-\gamma)N$ $\beta_{\mathcal{R},in}(1-\gamma)N$
12.5 MB/second 5910 73.88 GB/second
32.4 TB/month 191,484 TB/month
Appendix H Analysis of validation simulation
This supplementary section explains why state-dependent checks consume the majority of the processing time as illustrated in §8.1. Figure 16 dissects the distribution of the CPU processing time.
Ethereum uses the MPT to maintain the blockchain state, which is a tree structure where a leaf node stores the value of persistent data (e.g., account balance), and all intermediate nodes in the path from the root node to the leaf node correspond to the key of the data’s value (e.g., account address). Fetching account nonces also requires multiple lookups in the tree. This traversal process is time-consuming for several reasons [26]. First, due to the large amount of state, the tree can become deep, increasing the number of nodes that must be traversed to find a particular piece of data. Second, disk access plays a significant role because the MPT structure often resides on disk. Slow disk accesses occur when the nodes required for a lookup are not in memory and must be fetched from disk.
Although the nonce check appears to be the most expensive operation, it is not because other state-based validations (e.g., balances) are cheap; they appear faster primarily because when fetching the nonce, the balances of the account are also cached in one look-up, minimizing additional overhead (e.g., trie.(*StateTrie).GetAccount in Figure 16). This suggests that omitting one check while retaining another (i.e., merely checking the nonce but not the balance) does not alter the validation process, thereby not contributing meaningfully to latency reduction.
Figure 16: Profiling execution time for transaction validation — analysis of $136,437$ Transactions from $1,000$ consecutive blocks starting at block $18,140,000$ using Golang’s “pprof.”
Appendix I Discussions on measuring latency on testnet
In §8.2, we measure the latency (confirmation time) between the regular and the modified node at the test net Holesky. This section discusses how we process our measurements and explains the limitations.
We observe that users send a large number of transactions in a short period (e.g., attacks and tests), particularly due to the absence of transaction fees in testnet. To eliminate the effect of outliers, we count the number of transactions for five-minute intervals and exclude the period when the number of observed transactions is more than 3 standard deviations above the mean—excluding 1.14% of our total observation time.
Even after excluding abnormal activities, the market conditions differ significantly over time. Figure 17 shows the median confirmation time for each day. The colors represent the type of node running. It is clear that the transaction inclusion time varies by a margin of a few seconds each day, which could offset the latency reduction benefits gained from skipping validations.
In addition, the connected peers differ as well. Our node connects to a different set of peers even with the same public key (around 15% overlap), indicating that testnet peers might change very frequently. Furthermore, the testnet is much less active than the mainnet (around 30%). If the node receives many transactions at once, it may not be able to process transactions simultaneously, which adds latency to the regular node’s processing time. Finally, while this experiment only evaluates a single modified node, deploying or replaying multiple nodes without validation could lead to a cumulative latency reduction.
Figure 17: Time inclusion for a modified/regular node on a daily basis (median): Red (modified) & Blue (regular)